Govan Limited expects to purchase 40 000 components during 2021 with a purchase price of R20 per unit, an ordering cost of R2 per order and a carrying cost of 20 percent of the unit purchase price.Use this information given below to calculate the number of orders that should be placed during 2021 using the most advantageous quantity to order each time.


1
Expert's answer
2021-05-06T13:15:44-0400

To determine the number of orders that should be placed during 2021 using the most advantageous quantity, we use the Economic Order Quantity (EOQ) formula.


EOQ = "\sqrt{\frac{2DS}{H} }"

Annual Demand (D) = 40,000

Set up cost/order costs (S) = R 2 per order

Holding costs/Carrying costs (H) = 20% x R20 = R4 per order

Substitute these figures in the above formula:

EOQ (Q) = "\sqrt{\frac{2DS}{H} } = \sqrt{\frac{2\times 40000\times 2}{4} } = \sqrt{160000 } = 200" components

Therefore, the number of orders that should be placed during 2021 = "\frac{D}{Q} = \frac{40000}{200} = 200" orders